# On the uniform convexity of the squared distance

### v1

@@ 1 Introduction
The proper generalization of uniform convexity to this $W$-hyperbolic setting was found by Leuştean [17, 18] in the form of $UCW$-hyperbolic spaces. Some of the properties of uniformly convex normed spaces were found to generalize naturally to $UCW$-hyperbolic spaces in those papers, as well as in [14]. Typical examples of non-normed $UCW$-hyperbolic spaces are provided by the class of CAT(0) spaces, the canonical nonlinear generalization of inner product spaces, as studied in metric geometry. For a recent example of a $UCW$-hyperbolic space which is neither a CAT(0) space, nor a convex subset of a normed space, see [21, Section 4].
All these classes of nonlinear spaces, with their basic definitions and properties, may be found in Section 2. In addition, we show that the fixed point sets of (asymptotically) nonexpansive mappings on $UCW$-hyperbolic spaces are convex in a more general, quantitative way, as inspired by proof mining.
@@ 2.1 Nonlinear spaces
[…]
@@
As remarked in [17, Proposition 2.6], CAT(0) spaces are $UCW$-hyperbolic spaces having as a modulus of uniform convexity the simple function $(r,\varepsilon)\mapsto\varepsilon^{2}/8$.
In a metric space $(X,d)$, a nonempty subset $S$ of $X$ is a Chebyshev set if, for any $x\in X$, there exists a unique $y\in S$ such that, for any $z\in S$, $d(x,y)\leq d(x,z)$ – this defines the projection mapping $P_{S}:X\to S$. By [18, Proposition 2.4], any closed, convex, nonempty subset of a complete $UCW$-hyperbolic space is a Chebyshev set. Also, in any metric space $(X,d)$, we say that a sequence $(x_{n})$ is Fejér monotone with respect to a subset $S$ of $X$ if, for all $n\in\mathbb{N}$ and $q\in S$, $d(x_{n+1},q)\leq d(x_{n},q)$.
@@ 2.2 Nonexpansive mappings
A self-mapping $T$ of a metric space $(X,d)$ is called nonexpansive if, for any $x$, $y\in X$, $d(Tx,Ty)\leq d(x,y)$ – it is immediate that, for such a mapping, $Fix(T)$ is closed. By [18, Corollary 3.7], if $X$ is a bounded complete nonempty $UCW$-hyperbolic space, then any nonexpansive self-mapping of $X$ has a fixed point.
The following proposition, a quantitative generalization of the convexity of the fixed point set as inspired by proof mining, uses a quantitative version of the argument used in the proof of the strict convexity of $UCW$-hyperbolic spaces [17, Proposition 5].
@@
Let $\eta:(0,\infty)\times(0,2]\to(0,\infty)$ be monotone and let $(X,d,W)$ be a $UCW$-hyperbolic space admitting $\eta$ as a modulus. Let $T:X\to X$ be nonexpansive, $x$, $y\in X$, $\varepsilon\geq 0$, $b>0$ with $d(x,y)\leq b$. Let $\lambda\in[0,1]$ and set $z:=(1-\lambda)x+\lambda y$. Put
$\delta:=\min\left(\frac{\varepsilon}{2},b,\frac{\varepsilon^{2}}{16b}\eta\left(2b,\frac{\varepsilon}{2b}\right)\right).$
@@
so we are only interested in cases where $\frac{\varepsilon}{2b}\leq 2$, so the issue of the second argument of $\eta$ being restricted to $(0,2]$ does not come up.
@@
Case II. $1-\lambda\leq\frac{\varepsilon}{4b}$. Similarly.
@@ 2.3 Asymptotically nonexpansive mappings
If $(\delta_{n})$ is a sequence of non-negative reals such that $\delta_{n}\to 0$ (sometimes, the stronger condition $\sum\delta_{n}<\infty$ will be imposed), then a self-mapping $T$ of a metric space $(X,d)$ is called asymptotically nonexpansive w.r.t. $(\delta_{n})$ if, for any $x$, $y\in X$ and $n\in\mathbb{N}$, $d(T^{n}x,T^{n}y)\leq(1+\delta_{n})d(x,y)$ – it is immediate that, for such a mapping, $Fix(T)$ is closed. By [14, Theorem 3.3], if $X$ is a bounded complete nonempty $UCW$-hyperbolic space, then any asymptotically nonexpansive self-mapping of $X$ has a fixed point.
@@
$\Theta_{b,\eta}(\varepsilon)$ $:=\frac{1}{2}\cdot\min\left(\frac{\varepsilon}{2},b,\frac{\varepsilon^{2}}{16b}\eta\left(2b,\frac{\varepsilon}{2b}\right)\right)$
@@
    Let $\varepsilon\geq 0$ and $n\geq\gamma_{b,\eta,u}(\varepsilon)$. Assume that $d(x,T^{n}x)\leq\Theta_{b,\eta}(\varepsilon)$ and that $d(y,T^{n}y)\leq\Theta_{b,\eta}(\varepsilon)$. Then $d(z,T^{n}z)\leq\varepsilon$.
@@
    Let $\varepsilon\geq 0$ and $B\geq 0$ be such that, for every $n$, $\delta_{n}\leq B$. Assume that $d(x,Tx)\leq\Omega_{b,\eta,u,B}(\varepsilon)$ and that $d(y,Ty)\leq\Omega_{b,\eta,u,B}(\varepsilon)$. Then $d(z,Tz)\leq\varepsilon$.
@@
    First of all, put
    $\delta:=\min\left(\frac{\varepsilon}{2},b,\frac{\varepsilon^{2}}{16b}\eta\left(2b,\frac{\varepsilon}{2b}\right)\right),$
    so $\Theta_{b,\eta}(\varepsilon)=\delta/2$ and $\gamma_{b,\eta,u}(\varepsilon)=u\left(\frac{\delta}{2b}\right)$.
    Now, note that
    $d(T^{n}z,x)\leq d(T^{n}z,T^{n}x)+\delta/2\leq(1+\delta_{n})d(z,x)+\delta/2\leq d(z,x)+\delta/2+\delta/2=\lambda d(x,y)+\delta$
    and, similarly,
    $d(T^{n}z,y)\leq(1-\lambda)d(x,y)+\delta,$
    so
    $d(z,T^{n}z)\leq d(z,x)+d(T^{n}z,x)\leq 2\lambda d(x,y)+\delta\leq 3b,$
    so we are only interested in cases where $\frac{\varepsilon}{2b}\leq 2$, so the issue of the second argument of $\eta$ being restricted to $(0,2]$ does not come up.
@@
    In this case, as before,
    $d(z,T^{n}z)\leq 2\lambda d(x,y)+\delta\leq\varepsilon.$
    Case II. $1-\lambda\leq\frac{\varepsilon}{4b}$. Similarly.
@@
    In this case,
    $d(z,T^{n}z)\leq 2\lambda d(x,y)+\delta\leq\varepsilon.$
@@
    We have that $r_{1}\leq 2b$, so
    and $r_{1}\geq\lambda d(x,y)>\frac{\varepsilon^{2}}{16b}$.
    Assume towards a contradiction that $d(z,T^{n}z)>\varepsilon\geq\frac{\varepsilon}{2b}\cdot r_{1}$. Since $d(z,x)\leq r_{1}$ and $d(T^{n}z,x)\leq r_{1}$, we have that
    $d\left(\frac{z+T^{n}z}{2},x\right)\leq\left(1-\eta\left(r_{1},\frac{\varepsilon}{2b}\right)\right)\cdot r_{1}\leq\left(1-\eta\left(2b,\frac{\varepsilon}{2b}\right)\right)\cdot r_{1}<r_{1}-\frac{\varepsilon^{2}}{16b}\eta\left(2b,\frac{\varepsilon}{2b}\right)\leq r_{1}-\delta.$
    Similarly, we obtain that
    $d\left(\frac{z+T^{n}z}{2},y\right)<r_{2}-\delta,$
    so
    $d(x,y)\leq d\left(\frac{z+T^{n}z}{2},x\right)+d\left(\frac{z+T^{n}z}{2},y\right)<r_{1}+r_{2}-2\delta=d(x,y),$
    a contradiction.
@@
    Set $n:=N_{b,\eta,u,B}(\varepsilon)$. Now,
    $d(x,T^{n}x)\leq\sum_{i\in[0,n)}d(T^{i}x,T^{i+1}x)\leq\sum_{i\in[0,n)}(1+\delta_{i})d(x,Tx)\leq\Omega_{b,\eta,u,B}(\varepsilon)\cdot n\cdot(1+B)\leq\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right).$
    Similarly, we get that
    $d(x,T^{n+1}x)\leq\Omega_{b,\eta,u,B}(\varepsilon)\cdot(n+1)\cdot(1+B)=\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right)$
    and also that $d(y,T^{n}y)\leq\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right)$ and $d(y,T^{n+1}y)\leq\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right)$.
    Since both $n$ and $n+1$ are greater or equal to $\gamma_{b,\eta,u}\left(\frac{\varepsilon}{2+B}\right)$, we get that $d(z,T^{n}z)\leq\frac{\varepsilon}{2+B}$ and $d(z,T^{n+1}z)\leq\frac{\varepsilon}{2+B}$. We have that
    $d(Tz,T^{n+1}z)\leq(1+\delta_{1})d(z,T^{n}z)\leq(1+B)\cdot\frac{\varepsilon}{2+B}$
    and, thus, that
    $d(z,Tz)\leq d(z,T^{n+1}z)+d(Tz,T^{n+1}z)\leq\frac{\varepsilon}{2+B}+(1+B)\cdot\frac{\varepsilon}{2+B}=\varepsilon,$
    so we are done.
@@ 2.4 On the uniform convexity of the squared distance
For uniformly convex normed spaces, this property essentially goes back to Zălinescu [34, Section 4] – see also [32, 33] for more variations on this theme, and see [15, Proposition 2.4] (an instantiation of the more general [3, Proposition 3.2]) for a formula for this modulus which is easily computable in terms of the modulus of uniform convexity.
@@
Recently, Pinto [20] extended the aforementioned analysis of Kohlenbach and the author of Reich’s theorem to ‘uniformly smooth hyperbolic spaces’ which are in addition $UCW$-hyperbolic. At the time of the writing of [20], it was not known whether, generally, $UCW$-hyperbolic spaces satisfy property $(G)$; however, Pinto identified in [20, Proposition 5.4] (see also [26, Proposition 2.5]) a weaker property of them, called property $(M)$ in [21, Definition 2.8], as being enough for the proof to go through.
@@ 3 Main result
The following is the generalization to $UCW$-hyperbolic spaces of [15, Proposition 2.4], which, as said earlier, is an instantiation of the more general [3, Proposition 3.2].
@@
For any suitable $\eta$, $r$, $\varepsilon$, put
$\psi_{\eta}(r,\varepsilon):=\min\left(\frac{\left(\min\left(\frac{\varepsilon}{2},\frac{\varepsilon^{2}}{40r}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right)\right)\right)^{2}}{4},\frac{\varepsilon^{2}}{32}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right)\right).$
Let $\eta:(0,\infty)\times(0,2]\to(0,\infty)$ be monotone and let $(X,d,W)$ be a $UCW$-hyperbolic space admitting $\eta$ as a modulus. Then $(X,d,W)$ has property $(G)$ with modulus $\psi_{\eta}$.
@@
so the issue of the second argument of $\eta$ being restricted to $(0,2]$ does not come up.
@@
$\gamma:=\frac{\varepsilon^{2}}{32}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right),\quad\delta:=\min\left(\frac{\varepsilon}{2},\frac{4\gamma}{5r}\right)=\min\left(\frac{\varepsilon}{2},\frac{\varepsilon^{2}}{40r}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right)\right),$
so that $\delta\leq r$ and $\psi_{\eta}(r,\varepsilon)=\min\left(\frac{\delta^{2}}{4},\gamma\right)$. Also, put $\alpha:=d(x,a)$ and $\beta:=d(y,a)$ and assume w.l.o.g. that $\alpha\leq\beta$. Assume towards a contradiction that
@@
$d\left(\frac{x+z}{2},\frac{x+y}{2}\right)\leq\frac{1}{2}d(z,y)<\frac{\delta}{2},$
@@
$d\left(\frac{x+y}{2},a\right)\leq d\left(\frac{x+z}{2},a\right)+d\left(\frac{x+z}{2},\frac{x+y}{2}\right)<d\left(\frac{x+z}{2},a\right)+\frac{\delta}{2}.$
@@
$<\left(d\left(\frac{x+z}{2},a\right)+\frac{\delta}{2}\right)^{2}$
$=d^{2}\left(\frac{x+z}{2},a\right)+2d\left(\frac{x+z}{2},a\right)\cdot\frac{\delta}{2}+\left(\frac{\delta}{2}\right)^{2}$
$=d^{2}\left(\frac{x+z}{2},a\right)+\frac{\delta}{2}\left(2d\left(\frac{x+z}{2},a\right)+\frac{\delta}{2}\right)$
$\leq d^{2}\left(\frac{x+z}{2},a\right)+\frac{\delta}{2}\left(2r+\frac{r}{2}\right)$
@@
$d^{2}\left((1-\lambda)x+\lambda y,a\right)\leq(1-\lambda)d^{2}(x,a)+\lambda d^{2}(y,a)-2\lambda(1-\lambda)\psi_{\eta}(r,\varepsilon).$
@@
Applying [17, Proposition 5], we obtain that $z=W(x,y,\lambda)$.
@@
$\leq(1-\lambda)d^{2}(x,a)+\lambda d^{2}(y,a)-2\lambda(1-\lambda)\psi_{\eta}(r,\varepsilon),$

### v2

@@ 1 Introduction
The proper generalization of uniform convexity to this $W$-hyperbolic setting was found by Leuştean [19, 20] in the form of $UCW$-hyperbolic spaces. Some of the properties of uniformly convex normed spaces were found to generalize naturally to $UCW$-hyperbolic spaces in those papers, as well as in [15]. Typical examples of non-normed $UCW$-hyperbolic spaces are provided by the class of CAT(0) spaces, the canonical nonlinear generalization of inner product spaces, as studied in metric geometry. For a recent example of a $UCW$-hyperbolic space which is neither a CAT(0) space, nor a convex subset of a normed space, see [23, Section 4].
All these classes of nonlinear spaces, with their basic definitions and properties, may be found in Section 2. In addition, we show that the fixed point sets of (asymptotically) nonexpansive mappings on $UCW$-hyperbolic spaces are convex in a more general, quantitative way, as inspired by proof mining. (Proof mining can frequently help in finding the most natural logical form for a given property.)
@@ 2.1 Nonlinear spaces
A subset $C$ of a $W$-hyperbolic space $(X,d,W)$ is called convex if, for any $x$, $y\in C$ and $\lambda\in[0,1]$, $W(x,y,\lambda)\in C$. If $(X,d,W$) is a $W$-hyperbolic space, $x$, $y\in X$ and $\lambda\in[0,1]$, we denote the point $W(x,y,\lambda)$ by $(1-\lambda)x+\lambda y$. We will, also, mainly write $\frac{x+y}{2}$ for $\frac{1}{2}x+\frac{1}{2}y$.
@@
As remarked in [19, Proposition 2.6], CAT(0) spaces are $UCW$-hyperbolic spaces having as a modulus of uniform convexity the simple function $(r,\varepsilon)\mapsto\varepsilon^{2}/8$.
In a metric space $(X,d)$, a nonempty subset $S$ of $X$ is a Chebyshev set if, for any $x\in X$, there exists a unique $y\in S$ such that, for any $z\in S$, $d(x,y)\leq d(x,z)$ – this defines the projection mapping $P_{S}:X\to S$. By [20, Proposition 2.4], any closed, convex, nonempty subset of a complete $UCW$-hyperbolic space is a Chebyshev set. Also, in any metric space $(X,d)$, we say that a sequence $(x_{n})$ is Fejér monotone with respect to a subset $S$ of $X$ if, for all $n\in\mathbb{N}$ and $q\in S$, $d(x_{n+1},q)\leq d(x_{n},q)$.
@@ 2.2 Nonexpansive mappings
A self-mapping $T$ of a metric space $(X,d)$ is called nonexpansive if, for any $x$, $y\in X$, $d(Tx,Ty)\leq d(x,y)$ – it is immediate that, for such a mapping, $Fix(T)$ is closed. By [20, Corollary 3.7], if $X$ is a bounded complete nonempty $UCW$-hyperbolic space, then any nonexpansive self-mapping of $X$ has a fixed point.
The following proposition, a quantitative generalization of the convexity of the fixed point set as inspired by proof mining, uses a quantitative version of the argument used in the proof of the strict convexity of $UCW$-hyperbolic spaces [19, Proposition 5]. The proposition is a significant nonlinear generalization of the argument in [13, Lemma 2.3].
@@
Let $\eta:(0,\infty)\times(0,2]\to(0,\infty)$ be monotone and let $(X,d,W)$ be a $UCW$-hyperbolic space admitting $\eta$ as a modulus. Let $T:X\to X$ be nonexpansive, $x$, $y\in X$, $\varepsilon>0$, $b>0$ with $d(x,y)\leq b$. Let $\lambda\in[0,1]$ and set $z:=(1-\lambda)x+\lambda y$. Put
$\delta:=\min\left(\frac{\varepsilon}{2},b,\frac{\varepsilon^{2}}{16b}\eta\left(2b,\min\left(\frac{\varepsilon}{2b},2\right)\right)\right).$
@@
i.e. we are only interested in cases where $\frac{\varepsilon}{2b}\leq 2$, where
$\min\left(\frac{\varepsilon}{2b},2\right)=\frac{\varepsilon}{2b}.$
@@
Case II. $1-\lambda\leq\frac{\varepsilon}{4b}$. Similarly, using that
$d(z,Tz)\leq d(z,y)+d(Tz,y)\leq 2(1-\lambda)d(x,y)+\delta\leq\varepsilon.$
@@ 2.3 Asymptotically nonexpansive mappings
If $(\delta_{n})$ is a sequence of non-negative reals such that $\delta_{n}\to 0$ (sometimes, the stronger condition $\sum\delta_{n}<\infty$ will be imposed), then a self-mapping $T$ of a metric space $(X,d)$ is called asymptotically nonexpansive w.r.t. $(\delta_{n})$ if, for any $x$, $y\in X$ and $n\in\mathbb{N}$, $d(T^{n}x,T^{n}y)\leq(1+\delta_{n})d(x,y)$ – it is immediate that, for such a mapping, $Fix(T)$ is closed. By [15, Theorem 3.3], if $X$ is a bounded complete nonempty $UCW$-hyperbolic space, then any asymptotically nonexpansive self-mapping of $X$ has a fixed point.
@@
$\Theta_{b,\eta}(\varepsilon)$ $:=\frac{1}{2}\cdot\min\left(\frac{\varepsilon}{2},b,\frac{\varepsilon^{2}}{16b}\eta\left(2b,\min\left(\frac{\varepsilon}{2b},2\right)\right)\right)$
@@
    Let $\varepsilon>0$ and $n\geq\gamma_{b,\eta,u}(\varepsilon)$. Assume that $d(x,T^{n}x)\leq\Theta_{b,\eta}(\varepsilon)$ and that $d(y,T^{n}y)\leq\Theta_{b,\eta}(\varepsilon)$. Then $d(z,T^{n}z)\leq\varepsilon$.
@@
    Let $\varepsilon>0$ and $B\geq 0$ be such that, for every $n$, $\delta_{n}\leq B$. Assume that $d(x,Tx)\leq\Omega_{b,\eta,u,B}(\varepsilon)$ and that $d(y,Ty)\leq\Omega_{b,\eta,u,B}(\varepsilon)$. Then $d(z,Tz)\leq\varepsilon$.
@@
    First of all, put δ:= min( ε2,b,ε216bη(2b,min(ε2b,2))), so $\Theta_{b,\eta}(\varepsilon)=\delta/2$ and $\gamma_{b,\eta,u}(\varepsilon)=u\left(\frac{\delta}{2b}\right)$.
    Now, note that (using in the third inequality that $n\geq u(\delta/2b)$) d(T^nz,x) ≤d(T^nz,T^nx) + δ/2 ≤(1+δ_n)d(z,x) + δ/2 ≤d(z,x) + δ/2 + δ/2 = λd(x,y) + δ and, similarly, d(T^nz,y) ≤(1-λ)d(x,y) + δ, so d(z,T^nz) ≤d(z,x) + d(T^nz,x) ≤2λd(x,y) + δ≤3b, i.e. we are only interested in cases where $\frac{\varepsilon}{2b}\leq 2$, where min(ε2b,2) = ε2b.
@@
    In this case, as before, d(z,T^nz) ≤2λd(x,y) + δε.
    Case II. $1-\lambda\leq\frac{\varepsilon}{4b}$. Similarly, using that d(z,T^nz) ≤d(z,y) + d(T^nz,y) ≤2(1-λ)d(x,y) + δε.
@@
    In this case, d(z,T^nz) ≤2λd(x,y) + δε.
@@
    We have that $r_{1}\leq 2b$, so η(r_1,ε2b) ≥η(2b,ε2b), and $r_{1}\geq\lambda d(x,y)>\frac{\varepsilon^{2}}{16b}$.
    Assume towards a contradiction that $d(z,T^{n}z)>\varepsilon\geq\frac{\varepsilon}{2b}\cdot r_{1}$. Since $d(z,x)\leq r_{1}$ and $d(T^{n}z,x)\leq r_{1}$, we have that d(z+Tnz2,x) ≤(1-η(r_1,ε2b)) ⋅r_1 ≤(1-η(2b,ε2b)) ⋅r_1 ¡ r_1 - ε216bη(2b,ε2b) ≤r_1 - δ.
    Similarly, we obtain that d(z+Tnz2,y) ¡ r_2 -δ, so d(x,y) ≤d(z+Tnz2,x) + d(z+Tnz2,y) ¡ r_1 + r_2 - 2δ= d(x,y), a contradiction.
@@
    Set $n:=N_{b,\eta,u,B}(\varepsilon)$. Now, d(x,T^nx) ≤∑_i ∈[0,n) d(T^ix,T^i+1x) ≤∑_i ∈[0,n)(1+δ_i)d(x,Tx) ≤Ω_b,η,u,B(ε) ⋅n ⋅(1+B) ≤Θ_b,η(ε2+B). Similarly, we get that d(x,T^n+1x) ≤Ω_b,η,u,B(ε) ⋅(n+1) ⋅(1+B) = Θ_b,η(ε2+B) and also that $d(y,T^{n}y)\leq\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right)$ and $d(y,T^{n+1}y)\leq\Theta_{b,\eta}\left(\frac{\varepsilon}{2+B}\right)$.
    Since both $n$ and $n+1$ are greater or equal to $\gamma_{b,\eta,u}\left(\frac{\varepsilon}{2+B}\right)$, we get, using (i), that $d(z,T^{n}z)\leq\frac{\varepsilon}{2+B}$ and $d(z,T^{n+1}z)\leq\frac{\varepsilon}{2+B}$. We have that d(Tz,T^n+1z) ≤(1+δ_1)d(z,T^nz) ≤(1+B) ⋅ε2+B and, thus, that d(z,Tz) ≤d(z,T^n+1z) + d(Tz,T^n+1z) ≤ε2+B + (1+B) ⋅ε2+B = ε, so we are done.
@@ 2.4 On the uniform convexity of the squared distance
For uniformly convex normed spaces, this property essentially goes back to Zălinescu [36, Section 4] – see also [34, 35] for more variations on this theme, and see [17, Proposition 2.4] (an instantiation of the more general [3, Proposition 3.2]) for a formula for this modulus which is easily computable in terms of the modulus of uniform convexity.
@@
Recently, Pinto [22] extended the aforementioned analysis of Kohlenbach and the author of Reich’s theorem to ‘uniformly smooth hyperbolic spaces’ which are in addition $UCW$-hyperbolic. At the time of the writing of [22], it was not known whether, generally, $UCW$-hyperbolic spaces satisfy property $(G)$; however, Pinto identified in [22, Proposition 5.4] (see also [28, Proposition 2.5]) a weaker property of them, called property $(M)$ in [23, Definition 2.8], as being enough for the proof to go through.
@@ 3 Main result
The following is the generalization to $UCW$-hyperbolic spaces of [17, Proposition 2.4], which, as said earlier, is an instantiation of the more general [3, Proposition 3.2].
@@
Let $\eta:(0,\infty)\times(0,2]\to(0,\infty)$ be monotone and let $(X,d,W)$ be a $UCW$-hyperbolic space admitting $\eta$ as a modulus. For any suitable $r$, $\varepsilon$, put
$\psi_{\eta}(r,\varepsilon):=\min\left(\frac{\left(\min\left(\frac{\varepsilon}{2},\frac{\varepsilon^{2}}{96r}\eta^{2}\left(r,\min\left(\frac{\varepsilon}{2r},2\right)\right)\right)\right)^{2}}{4},\frac{\varepsilon^{2}}{32}\eta^{2}\left(r,\min\left(\frac{\varepsilon}{2r},2\right)\right)\right).$
Then $(X,d,W)$ has property $(G)$ with modulus $\psi_{\eta}$.
@@
i.e. we are only interested in cases where $\frac{\varepsilon}{2r}\leq 2$, where
$\min\left(\frac{\varepsilon}{2r},2\right)=\frac{\varepsilon}{2r}.$
@@
$\gamma:=\frac{\varepsilon^{2}}{32}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right),\quad\delta:=\min\left(\frac{\varepsilon}{2},\frac{\gamma}{3r}\right)=\min\left(\frac{\varepsilon}{2},\frac{\varepsilon^{2}}{96r}\eta^{2}\left(r,\frac{\varepsilon}{2r}\right)\right),$
so that $\delta\leq\varepsilon/2\leq r$ and $\psi_{\eta}(r,\varepsilon)=\min\left(\frac{\delta^{2}}{4},\gamma\right)$. Also, put $\alpha:=d(x,a)$ and $\beta:=d(y,a)$ and assume w.l.o.g. that $\alpha\leq\beta$. Assume towards a contradiction that
@@
$d\left(\frac{x+z}{2},\frac{x+y}{2}\right)\leq\frac{1}{2}d(z,y)\leq d(z,y)<\delta,$
@@
$d\left(\frac{x+y}{2},a\right)\leq d\left(\frac{x+z}{2},a\right)+d\left(\frac{x+z}{2},\frac{x+y}{2}\right)<d\left(\frac{x+z}{2},a\right)+\delta.$
@@
$<\left(d\left(\frac{x+z}{2},a\right)+\delta\right)^{2}$
$=d^{2}\left(\frac{x+z}{2},a\right)+2d\left(\frac{x+z}{2},a\right)\cdot\delta+\delta^{2}$
$=d^{2}\left(\frac{x+z}{2},a\right)+\delta\left(2d\left(\frac{x+z}{2},a\right)+\delta\right)$
$\leq d^{2}\left(\frac{x+z}{2},a\right)+\delta\left(2r+r\right)$
@@
Remark 3.2.
An inspection of the proof shows that the only use of the axiom $(W4)$ may be replaced by that of the axiom $(W5)$, as introduced in [16], which states that, for all $x$, $y$, $z\in X$ and $\lambda\in[0,1]$,
$d(W(x,z,\lambda),W(y,z,\lambda))\leq d(x,y).$
@@
$d^{2}\left((1-\lambda)x+\lambda y,a\right)$ $\leq(1-\lambda)d^{2}(x,a)+\lambda d^{2}(y,a)-2\min(\lambda,1-\lambda)\psi_{\eta}(r,\varepsilon)$
$\leq(1-\lambda)d^{2}(x,a)+\lambda d^{2}(y,a)-2\lambda(1-\lambda)\psi_{\eta}(r,\varepsilon).$
@@
Applying [19, Proposition 5], we obtain that $z=W(x,y,\lambda)$.
@@
$\leq(1-\lambda)d^{2}(x,a)+\lambda d^{2}(y,a)-2\min(\lambda,1-\lambda)\psi_{\eta}(r,\varepsilon),$